OpenStudy (anonymous):

Which of the following is 0.212121... written as a fraction? 29 733 337 21100

jimthompson5910 (jim_thompson5910):

This page offers a good example http://www.mathrecreation.com/2009/03/decimals-to-fractions.html

OpenStudy (anonymous):

@PRAETORIAN.10 @arabpride @$w3G_Godd @surjithayer

OpenStudy (anonymous):

let x=0.212121......(1) 10x=2.12121.... 100 x=21.2121....(2) (2)-(1) gives 100 x-x=21 x=21/99=7/33

OpenStudy (anonymous):

so answer is?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

obviously733